OpenStudy (anonymous):

(Chemistry question) In the reaction C6H5OH + O2 --> CO2 + H2O, how many grams of phenol are required to react with .200 moles of oxygen? (equation isn't balanced yet)

OpenStudy (matt101):

First balance the equation: 2 C6H5OH + 15 O2 --> 12 CO2 + 6 H2O Phenol (C6H5OH) reacts with O2 in a 2:15 ratio. This means: moles phenol 2 ------------ = --- moles O2 15 moles phenol = (2/15)*0.200 = 0.0267 Molar mass of phenol is 94 g/mol, so the mass of phenol would be 0.0267*94 = 2.5 g